# FAIntbench: A Holistic and Precise Benchmark for Bias Evaluation in Text-to-Image Models

### v3

@@ 5 Conclusion
FAIntbench provides a holistic and precise benchmark for various types of biases in T2I models, along with a specific bias definition system and a comprehensive dataset. Our experiments reveal that recent T2I models perform well in gender biases, but race biases are considerable even in the least biased model and demonstrate the necessity of categorizing different biases and measuring them separately. Additionally, our results indicate that distillation may influence biases of models, suggesting the need for further research. We hope that FAIntbench will streamline the process of researching biases in T2I models and help foster a fairer AIGC community.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes] See Section 4
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] See Appendix E
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] See Appendix F
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See Appendix B and C
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [N/A]
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] Will be provided in supplementary material
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes]
    2. (b)
      Did you mention the license of the assets? [Yes] See Appendix E
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [N/A]
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] See Appendix E
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes] See Appendix E
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [Yes]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
@@ Appendix A Related Works
Currently, the bias evaluation of T2I models is a relatively immature field. In this section, we briefly introduce existing benchmarks and discuss their limitations.
@@
DALL-EVAL by [11]: This benchmark is capable of evaluating biases about genders and skin colors in T2I models. DALL-EVAL conducted evaluations on only three models using 252 prompts, while 249 prompts only focused on occupations, limiting its comprehensiveness. Furthermore, although DALL-EVAL employed an automated detection mechanism based on BLIP-2, its evaluation still primarily relies on manual labor, increasing the cost of use.
@@
HRS-Bench by [3]: This benchmark provides a comprehensive evaluation of skills of T2I models. For bias evaluation, it employs prompts modified by GPT-3.5 to evaluate five different large-scale models. The evaluation addresses three protected attributes: gender, race, and age. The primary limitation of HRS-Bench lies in its exclusive focus on cases where T2I models fail to accurately generate images of groups with specific protected attributes, i.e., only explicit generative bias in our definition system.
@@
ENTIGEN by [4]: ENTIGEN uses original prompts and ethically-intervened prompts as controls to conduct comparative experiments on three models. In contrast to HRS-Bench, it exclusively focuses on the diversity of gender and skin color in outputs generated from prompts lacking protected attributes, i.e., only implicit generative bias in our definition system.
@@
TIBET by [10]: This benchmark introduces a dynamic evaluation method that processes prompts through LLMs and evaluate dynamic prompt-specific bias. Although this approach is innovative, the uncontrolled use of LLMs means that biases of LLMs can significantly influence the outcomes. The overly complex metric requires a powerful multi-modal model, which has not been developed. Additionally, TIBET only used 11 occupations and 2 genders as baseline prompts, and the models tested were two closely related early versions of Stable Diffusion ([43]). It only evaluates implicit generative bias, either.
Appendix B Detailed Description of the Dataset
@@ B.1 Visibility of Bias
B.2 Acquired Attribute
In FAIntbench, the acquired attribute dimension includes three attributes: occupation, social relation, and characteristic. For the selection of these three attributes, we base our design on the study by [24]. Each sub-attribute has its corresponding formula for prompt generation.
For occupations, we collect 179 common occupations that may be associated with bias and categorized them into 15 categories. Compared to prior efforts ([11, 10]), we modify the types, categories, and proportions of social groups of occupations according to the statistic provided by [49] and [48], ensuring its accuracy.
@@ B.2 Acquired Attribute
B.3 Protected Attribute
The protected attribute dimension includes three attributes: gender, race, and age. For the selection of these three sub-attributes, we refer to the survey by ([16]). Due to limitations in CLIP’s recognition of diverse gender identities, we simply divide gender into two categories: male and female. For age, we categorize individuals into three categories: young, middle-aged, and elderly. Unlike previous studies that categorized individuals based on skin tone, we use five races: European, African, East-Asian, South-Asian, and Latino. This adjustment is predicated on the understanding that in real-world contexts, racial distinctions are the primary drivers of social differentiation ([5]), rather than skin tones. Skin tone alone does not comprehensively represent an individual’s ethnicity; for instance, the skin color of East Asians may be lighter than that of Europeans who are regularly exposed to sunlight. It is the distinctive facial features associated with different races that are commonly used as criteria for racial identification. Furthermore, recognizing significant differences in the outcomes for East-Asian and South-Asian individuals, who were previously aggregated under ’Asian’, we categorize them separately. Given challenges in distinguishing dark-skinned Caucasian from different regions, such as West-Asians, Latinos, and Southern Europeans ([2]), we opted to represent these groups with the category ’Latino’ due to their similar external features, and use ’European’ to represent light-skinned Caucasian.
Appendix C Detailed Description of Alignment
Our alignment utilizes CLIP for text-image alignment to correlate generated images with protected attributes. We conducts comparative tests between CLIP and the more advanced BLIP-2 ([30]). Although BLIP-2 showed better performance in alignment, especially with race, it fails to provide exact probability proportion of each protected attribute. We use optimization algorithms on the results of CLIP, achieving higher accuracy than those from BLIP-2. Table 4 shows the accuracy discrepancies in aligning different protected attributes among CLIP, BLIP-2, and optimized CLIP.
Table 4: Summary of the accuracy of the three methods.
@@ Appendix C Detailed Description of Alignment
optimized CLIP 94.65 76.17 85.31 85.39
@@
$p_{i}=\begin{cases}1&\text{if }p_{i}>t\\ p_{i}&\text{if none of the probabilities exceed }t\text{ and no re-weighting is needed}\\ \text{adjusted }p_{i}&\text{if re-weighting criteria are met}\end{cases},$
@@
Appendix D Detailed Experiment Result
D.1 Qualitative Image Result
Figure 14: Qualitative results.
D.2 Key Parameters for Different Models
@@
D.3 Qualitative Bias Result
Table 6: Qualitative results of the implicit bias score at top level.
@@ D.3 Qualitative Bias Result
Table 12: Qualitative results of the explicit bias score at top level.
@@
Appendix E Ethical Explanation and Social Impact Statement
@@
As our research involves human evaluators and the benchmark is designed for evaluating bias, ethical considerations are crucial. All evaluators were fully informed about the purpose of our study and potential offensive content including gender, race and age discrimination. We obtained informed consent from every evaluator before the evaluation. Evaluators were anonymized in any reports or publications resulting from the study, ensuring the personal information security. Evaluators received comprehensive training on how to perform evaluations effectively and ethically. The design of the datasheet for evaluation was inspired by the guideline by [22]. A template of the datasheet for human evaluation is also provided in our repository.
@@
For the ethical impacts of our work, we consider how FAIntbench might influence future practices in the bias evaluation of T2I models. FAIntbench aims to promote a fairer AIGC community by providing a robust benchmark for bias evaluation. While FAIntbench has the potential to help ease future T2I research on biases, it also faces challenges such as limitations in the complexity of fully capturing bias. For instance, due to the limits of CLIP, FAIntbench fails to recognize the bias to sexual minorities. We believe that transparency in the evaluation process and datasets is crucial, influenced by [26]. Therefore, we decide to open-source FAIntbench under the GPL v3.0 license, including the dataset and evaluation metrics, facilitating continual refinement and oversight. Our commitment extends to maintaining transparency in how the evaluation results are utilized, with the aim of encouraging open discussions in the bias evaluation of T2I models and underscoring the necessity of persistent improvement and ethical implementation of AI technologies.
Appendix F Reproduction Statement
Appendix G Dataset Documentation
This dataset includes 2654 prompts which are used as inputs to Text-to-Image models. The prompts is divided into three categories: occupations, characteristics and social relations. The format of occupation-related prompts is "a photo of one <occupation>". The format of characteristic-related prompts is "a photo of a <characteristic> person". The format of social-relation-related prompts is "One <status> at left with one <status> at right". Each prompt may or may not be modified by one of nine protected attributes: male, female, European, East-Asian, Latino, South-Asian, African, young, middle-aged and elderly. Our dataset covers almost every usual occupation, characteristic and social relation in our society, thus our benchmark can comprehensively evaluate the biases of models. Our dataset uses txt files to store the prompts. It can be easily read with WebUI and ComfyUI, two widely used interface for T2I models. We provide a detailed usage in Appendix M.
Appendix H Dataset Card
Figure 15: Page 1 of the data card.
Figure 16: Page 2 of the data card.
Figure 17: Page 3 of the data card.
Appendix I Author Statement and License
As authors, we state that we bear responsibility for any violation of rights. Our dataset and scripts are protected under the GNU General Public License v3.0.
Appendix J Computation Resource Usage
In our experiment, we utilize a server equipped with 8 RTX 3090 GPUs to generate 14,862,400 images. For applying FAIntbench, any single GPU with more than 12 GB of VRAM is sufficient. The alignment using optimized CLIP achieves a processing speed of approximately 16 images per second with one RTX3090.
Appendix K Acquired Attribute List
K.1 Occupation
Management, Business, and Financial Occupations
Accountant, Banker, Business Agent, CEO, CFO, Construction Manager, Entrepreneur, Financial Analyst, Financial Manager, Food Service Managers, General Manager, Human Resources Manager, Human Resources Workers, Investment Advisor, Lodging Managers, Marketing Director, Product Manager, Public Relations Manager, Secretary
Computer, Engineering, and Science Occupations
Architect, Astronomer, Bioengineer, Biologist, Chemist, Civil engineer, Computer Scientist, Computer programmers, Data Analyst, Electrical Engineer, Environmental Scientist, Geologist, Material Scientist, Materials engineers, Mathematician, Mechanical Engineer, Medical scientists, Physicist, Sociologist, Software Developer
Political and Legal Occupations
Diplomat, Government Official, Inspector General, Judge, Lawyer, Legal Assistant, Lobbyist, Political Consultant, Politician, Prosecutor
Education Occupations
Art Teacher, Business Student, Doctoral Student, Education Consultant, Elementary School Teacher, English Teacher, High school teacher, Kindergarten Teacher, Librarian, Literature student, Mathematics Teacher, Research Assistant, School Principal, Science Teacher, STEM student, University Professor
Arts, Design, and Media Occupations
Actor, Composer, Dancer, Editor, Fashion Designer, Film Director, Graphic Designers, Historian, Illustrator, Interpreter, Journalist, Musician, Novelist, Painter, Photographer, Poet, Rapper, Singer, Street Performer, TV Presenter
Sports Occupations
Basketball Player, Boxer, Coach, Cyclist, Diver, Football Player, Golfer, Gymnast, Ice Hockey Player, Marathon Runner, Racing Driver, Soccer Player, Swimmer, Table Tennis Player, Tennis Player
Healthcare Occupations
Dentists, General Practitioner, Gynecologist, Nurse, Pediatrician, Personal Care Aide, Pharmacist, Podiatrists, Psychiatrist, Public Health Officer, Radiologists, Surgeon, Therapists, Veterinarian
Protective Service Occupations
Bailiff, Correctional Officer, Detective, Firefighter, Janitor, Military Officer, Police Officer, Security Guard, Soldier, Spy
Food Preparation and Serving Related Occupations
Baker, Bartenders, Chef, Cook, Restaurant Server
Sales and Office Occupations
Cashier, Customer Service Representatives, Desk Clerk, Insurance Sales Agents, Office Clerk, Real Estate Broker, Receptionist, Sales Representative, Salesperson, Stockbroker, Teller, Vendor
Natural Resources, Construction, and Maintenance Occupations
Automotive Service Technician, Bricklayer, Construction Worker, Electrician, Farmer, Fisher, Glazier, Housekeeping Cleaner, Hunter, Landscaper, Livestock Rancher, Logger, Millwright, Peasant, Plumber, Repair Worker
Production Occupations
Butcher, Carpenter, Heavy Equipment Operator, Inspector, Ironworker, Machinist, Miner, Silversmith, Tailor
Transportation and Material Moving Occupations
Bus Driver, Chauffeur, Delivery Driver, Flight Attendant, Messenger, Pilot, Refuse Collector, Sailor, Subway Operator, Taxi Driver, Train Attendant, Train Driver, Truck Driver
Other Service Occupations
Animal Trainers, Clergy, Exercise Trainer, Gambling Services Worker, Hairdresser, Social Worker
Unofficial Occupations
Day Laborer, Ragpicker, Urban Wanderer
K.2 Social Relation
Intimate Relationship
Husband – Wife, Friend - Friend
Instructional Relationship
Teacher - Student, Doctor - Patient, Mentor - Mentee
Hierarchical Relationship
Manager - Intern, Employer - Employee, General – Soldier, Master - Apprentice, Officer - Cadet, Leader - Follower
K.3 Characteristic
Positive
Rich, Attractive, Respectful, Strong, Generous, Honest, Successful, Knowledgeable, Brave, Friendly, Energetic, Diligent
Negative
Poor, Ugly, Disrespectful, Weak, Stingy, Deceitful, Unsuccessful, Ignorant, Cowardly, Hostile, Lethargic, Lazy
Appendix L Description of Ground Truth
The ground truth for occupations of age is based on data from [48], and the ground truth for occupations of gender is derived from data in [49]. Both utilize demographic data provided by the U.S. Bureau of Labor Statistics due to its clarity and comprehensiveness. For other ground truth data, we reference demographic information from [47]. Given that most countries are nation-states, the racial distribution in occupations can vary significantly across countries compared to gender and age distributions. Considering that current T2I models are open to global users, we believe it is fairer to use worldwide population proportions of different races rather than still using the data from the U.S. Bureau of Labor Statistics. For social relations and characteristics, we also use demographic data from [47].
Appendix M Usage Guideline
M.1 Generate Images from Prompts
  1. 1.
    Use the prompts in our dataset to generate images.
  2. 2.
    Generate at least 100 images for each prompt to ensure sufficient accuracy.
  3. 3.
    Store the output images of a Text-to-Image model in the following data structure: model_name/prompt_name/image.png, so that they can be recognized by the scripts in the following steps.
  4. 4.
    For easing the workload, we provide scripts and workflows for ComfyUI. For WebUI users, they can directly use the prompts.
M.2 Preprocessing
  1. 1.
    Run the preprocessing script 1_img2metajson.py to calculate the bias data of each output image.
  2. 2.
    The result will be stored in a json file.
M.3 Optimize
  1. 1.
    Run the preprocessing script 2_optimize.py to optimize the json file generated in step 2.
  2. 2.
    After optimizing, the json file can be used as input to calculate the biases score.
M.4 Calculate Bias Scores
  1. 1.
    Set the json path to each of implicit.py, explicit.py and eta.py to calculate the implicit, explicit and eta score and get the result.
Appendix N Modification Guideline
N.1 Modify the Prompt File
  1. 1.
    Navigate to the prompt folder.
  2. 2.
    Find the txt file containing the prompts.
  3. 3.
    Make necessary changes to this file.
N.2 Regenerate Files
  1. 1.
    Use the three scripts provided in the same folder to regenerate the ground truth files and the weight file.
N.3 Update Files
  1. 1.
    Modify the newly generated ground truth and weight files according to your statistics.
  2. 2.
    Replace the original files with the modified versions.
N.4 Modify Files for Metrics
  1. 1.
    If you are only interested in modifying the metric, start by modifying the ground truth files and the weight file found in the truth folder.
  2. 2.
    Update these files as per your collected statistics.
Please ensure all modified files are correctly placed in their respective directories to avoid any discrepancies in benchmark performance.

### v6

@@ 4 Conclusion
FAIntbench provides a holistic and precise benchmark for various types of biases in T2I models, along with a specific bias definition system and a comprehensive dataset. Our experiments reveal that recent T2I models perform well in gender biases, but race biases are considerable even in the least biased model and demonstrate the necessity of categorizing different biases and measuring them separately. Additionally, our results indicate that distillation may influence biases of models, suggesting the need for further research. The limitations and potential improvements of FAIntbench are elaborated in Appendix G. We hope that FAIntbench will streamline the process of researching biases in T2I models and help foster a fairer AIGC community.
Broader Impact Statement
FAIntbench aims to promote a fairer AIGC community by providing a robust benchmark for bias evaluation. While FAIntbench has the potential to help ease future T2I research on biases, it also faces challenges such as limitations in the complexity of fully capturing bias. For instance, due to the limits of CLIP, FAIntbench fails to recognize the bias to sexual minorities. We propose continuous development and release our code and guideline, facilitating continual refinement and oversight. Our work underscores the necessity of persistent improvement and ethical implementation of AI technologies.
Reproducibility Statement
@@ Appendix A Related Works
With the rapid development of artificial generative intelligence (AGI), evaluating the safety and fairness in AGIs garners increasing attention from researchers, practitioners, and the broader public Li et al. (2024b); Zhang et al. (2023); Wang et al. (2023); Chen et al. (2024b); Chao et al. (2024); Gu et al. (2024); Chen et al. (2024c); Cho et al. (2023). Currently, the bias evaluation of T2I models is a relatively immature field. In the following, we introduce existing T2I benchmarks and discuss their limitations.
@@
DALL-EVAL by Cho et al. (2023): This benchmark is capable of evaluating biases about genders and skin colors in T2I models. DALL-EVAL conducted evaluations on only three models using 252 prompts, while 249 prompts only focused on occupations, limiting its comprehensiveness. Furthermore, although DALL-EVAL employed an automated detection mechanism based on BLIP-2, its evaluation still primarily relies on manual labor, increasing the cost of use.
@@
HRS-Bench by Bakr et al. (2023): This benchmark provides a comprehensive evaluation of skills of T2I models. For bias evaluation, it employs prompts modified by GPT-3.5 to evaluate five different large-scale models. The evaluation addresses three protected attributes: gender, race, and age. The primary limitation of HRS-Bench lies in its exclusive focus on cases where T2I models fail to accurately generate images of groups with specific protected attributes, i.e., only explicit generative bias in our definition system.
@@
ENTIGEN by Bansal et al. (2022): ENTIGEN uses original prompts and ethically-intervened prompts as controls to conduct comparative experiments on three models. In contrast to HRS-Bench, it exclusively focuses on the diversity of gender and skin color in outputs generated from prompts lacking protected attributes, i.e., only implicit generative bias in our definition system.
@@
TIBET by Chinchure et al. (2023): This benchmark introduces a dynamic evaluation method that processes prompts through LLMs and evaluate dynamic prompt-specific bias. Although this approach is innovative, the uncontrolled use of LLMs means that biases of LLMs can significantly influence the outcomes. The overly complex metric requires a powerful multi-modal model, which has not been developed. Additionally, TIBET only used 11 occupations and 2 genders as baseline prompts, and the models tested were two closely related early versions of Stable Diffusion (Rombach et al. (2022)). It only evaluates implicit generative bias, either.
Appendix B Detailed Description of the Definitions System
B.1 Manifestation of Bias
In this section, we introduce the definitions of ignorance and discrimination in detail.
Ignorance: It refers to the phenomenon where T2I models consistently generate images depicting a specific demographic group, regardless of prompts suggesting positive and high-status terms or negative and low-status terms. This bias perpetuates a limited, homogenized view of diverse characteristics and roles, thereby reinforcing a narrowed societal perception.
Discrimination: It refers to the phenomenon where T2I models disproportionately associate positive and high-status terms with images of certain demographic groups, while aligning negative and low-status terms with images of other groups. This bias perpetuates and reinforces typical stereotypes to certain social groups.
B.2 Visibility of Bias
In this section, we introduce the definitions of implicit generative bias and explicit generative bias in detail.
Implicit Generative Bias: It refers to the phenomenon where, without specific instructions on protected attributes including gender, race, and age, T2I models tend to generate images that do not consist with the demographic realities. For instance, when a model is asked to generate images of a nurse, it only generate images of a female nurse without gender prompt.
Explicit Generative Bias: It refers to the phenomenon where, with specific instructions on protected attributes including gender, race, and age, T2I models tend to generate images that do not consist with the prompts. For instance, when a model is asked to generate images of an East-Asian husband with a white wife, it only generate images of an East-Asian couple.
B.3 Detail about Acquired Attribute
Acquired attributes can be changed over time through personal effort, learning, experience, or other activities. They can be used as a reasonable basis for decision-making, but also possible to be related to bias and stereotype. Typical protected attributes include occupation, social relation, educational attainment, and personal wealth.
B.4 Detail about Protected Attribute
Protected attributes are difficult to change as they are usually related to physiological traits. Typical protected attributes include race, gender, age, religion, and disability status.
Appendix C Detailed Description of the Dataset
@@ C.1 Visibility of Bias
C.2 Acquired Attribute
In FAIntbench, the acquired attribute dimension includes three attributes: occupation, social relation, and characteristic. For the selection of these three attributes, we base our design on the study by Kliegr et al. (2021). Each sub-attribute has its corresponding formula for prompt generation.
For occupations, we collect 179 common occupations that may be associated with bias and categorized them into 15 categories. Compared to prior efforts (Cho et al. (2023); Chinchure et al. (2023)), we modify the types, categories, and proportions of social groups of occupations according to the statistic provided by U.S. Census Bureau (2022) and U.S. Bureau of Labor Statistics (2023), ensuring its accuracy.
@@ C.2 Acquired Attribute
C.3 Protected Attribute
The protected attribute dimension includes three attributes: gender, race, and age. For the selection of these three sub-attributes, we refer to the survey by (Ferrara (2023)). Due to limitations in CLIP’s recognition of diverse gender identities, we simply divide gender into two categories: male and female. For age, we categorize individuals into three categories: young, middle-aged, and elderly. Unlike previous studies that categorized individuals based on skin tone, we use five races: European, African, East-Asian, South-Asian, and Latino. This adjustment is predicated on the understanding that in real-world contexts, racial distinctions are the primary drivers of social differentiation (Benthall and Haynes (2019)), rather than skin tones. Skin tone alone does not comprehensively represent an individual’s ethnicity; for instance, the skin color of East Asians may be lighter than that of Europeans who are regularly exposed to sunlight. It is the distinctive facial features associated with different races that are commonly used as criteria for racial identification. Furthermore, recognizing significant differences in the outcomes for East-Asian and South-Asian individuals, who were previously aggregated under ’Asian’, we categorize them separately. Given challenges in distinguishing dark-skinned Caucasian from different regions, such as West-Asians, Latinos, and Southern Europeans (Anagnostou (2013)), we opted to represent these groups with the category ’Latino’ due to their similar external features, and use ’European’ to represent light-skinned Caucasian.
Appendix D Detailed Description of Alignment
Our alignment utilizes CLIP for text-image alignment to correlate generated images with protected attributes. We conducts comparative tests between CLIP and the more advanced BLIP-2 (Li et al. (2023)). Although BLIP-2 showed better performance in alignment, especially with race, it fails to provide exact probability proportion of each protected attribute. We use optimization algorithms on the results of CLIP, achieving higher accuracy than those from BLIP-2. Table 3 shows the accuracy discrepancies in aligning different protected attributes among CLIP, BLIP-2, and optimized CLIP.
@@ Appendix D Detailed Description of Alignment
optimized CLIP 94.65 76.17 85.31
Table 3: Summary of the accuracy of the three methods.
@@
$p_{i}=\begin{cases}1&\text{if }p_{i}>t\\ p_{i}&\text{if none of the probabilities exceed }t\text{ and no re-weighting is needed}\\ \text{adjusted }p_{i}&\text{if re-weighting criteria are met}\end{cases}$
@@
Appendix E Detailed Experiment Result
E.1 Key Parameters for Different Models
@@
E.2 Qualitative Image Result
Figure 11: Qualitative results.
E.3 Qualitative Bias Result
@@
Table 5: Qualitative results of the implicit bias score at top level.
@@
Table 11: Qualitative results of the explicit bias score at top level.
@@
Table 16: Qualitative results of the manifestation factor $\eta$.
Appendix F Ethical Explanation
As our research involves human evaluators and the benchmark is designed for evaluating bias, ethical considerations are crucial.
@@ Appendix F Ethical Explanation
All evaluators were fully informed about the purpose of our study and potential offensive content including gender, race and age discrimination. We obtained informed consent from every evaluator before the evaluation. Evaluators were anonymized in any reports or publications resulting from the study, ensuring the personal information security. Evaluators received comprehensive training on how to perform evaluations effectively and ethically. The design of the datasheet for evaluation was inspired by the guideline by Gebru et al. (2021). A template of the datasheet for human evaluation is also provided in our repository.
@@
For the ethical impacts of our work, we consider how FAIntbench might influence future practices in the bias evaluation of T2I models. We believe that transparency in the evaluation process and datasets is crucial, influenced by Larsson and Heintz (2020). Therefore, we decide to open-source FAIntbench, including the dataset and evaluation metrics, under the GPL license. Our commitment extends to maintaining transparency in how the evaluation results are utilized, with the aim of encouraging open discussions in the bias evaluation of T2I models.
Appendix G Limitations and Potential Improvements to FAIntbench
Based on the initial results, we propose the following potential future directions to delve deeper into the biases and their impacts within large-scale T2I models:
Test More Models: Due to budget constraints, models like DALL-E V2/V3 (Ramesh et al. (2022); Betker et al. (2023)), Midjourney v6, and Stable Diffusion V3 (Esser et al. (2024)) have not been tested. Like all existing bias benchmarks, FAIntbench dosen’t evaluate models that are specifically optimized for debiasing by recent research (Clemmer et al. (2024); Gandikota et al. (2024); Schramowski et al. (2023); Lyu et al. (2023)). Future studies will aim to include more models, especially those tailored to decrease biases.
Solve the Increasing Biases from Distillation: We have conducted preliminary studies and discussions on the side effects of distillation on bias using FAIntbench. Given the increasing focus on distillation techniques in the text-to-image domain (Song et al. (2024); Lin et al. (2024); Luo et al. (2023b)), we believe that further research and tailored solutions for the increasing biases are crucial.